OpenStudy (anonymous):

Please help what is the solution to the system of equations? 5x - 3y = -9 2x - 5y = 4 ( , )

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@Joel_the_boss

JoelTheBoss (joel_the_boss):

Do you know anything about this particular question? Is there a specific step you need help with?

OpenStudy (anonymous):

i need to know how to change it so y is alone on one side

OpenStudy (anonymous):

or the answer :)

JoelTheBoss (joel_the_boss):

Ok first you should add 3y to both sides \[5x=-9+3y\] Next just regroup the terms,\[5x=3y-9\] Next step would be to divide 5 on both sides, \[x=\frac{ 3y-9 }{ 5}\] After this Just factor out the common term 3, \[x=\frac{ 3(y-3) }{ 5}\] Did that help? Are you beginning to understand?

OpenStudy (anonymous):

yes kind of

OpenStudy (anonymous):

but it dosent help me with the answer

OpenStudy (triciaal):

there are 4 different ways to solve simultaneous equations but the goal is to find the set of values that will satisfy both equations at the same time.

JoelTheBoss (joel_the_boss):

Oh wait hehe, don't do that last step :P, I forgot its a system of equations, \[x=\frac{ 3y }{ 5}-\frac{ 9 }{ 5}\]That's x. :D Now, that you know this, plug that in for x to get y.

OpenStudy (anonymous):

lol yeah i was getting confused

OpenStudy (triciaal):

for this one let's use elimination we want the coefficients of one variable to be equal and opposite so they cancel out. this will give one equation with one unknown. solve for this variable then substitute in either of the original to find the value of the other variable

OpenStudy (triciaal):

multiply the first equation by 2 and the 2nd equation by -5 what do you get?

OpenStudy (anonymous):

can you show me im a little confused

OpenStudy (triciaal):

which is the first equation

OpenStudy (anonymous):

5x - 3y = -9

OpenStudy (triciaal):

multiply each term by 2

OpenStudy (anonymous):

10x - 6y = -18

OpenStudy (triciaal):

lets name this equation 3

OpenStudy (triciaal):

now multiply equation 2 by -5 and name that equation 4

OpenStudy (anonymous):

2x - 5y = 4 -10 - 25 = -20

OpenStudy (anonymous):

-10x - 25y = -20

OpenStudy (triciaal):

good do you see that the coefficients of x are equal and opposite for equations 3 and 4 add the equations together and solve for y

OpenStudy (anonymous):

how do i add them?

OpenStudy (anonymous):

x - 19 = -2

OpenStudy (anonymous):

is that right?

OpenStudy (triciaal):